# Interactive Multiscale Modeling to Bridge Atomic Properties and Electrochemical Performance in Li-CO2 Battery Design

### v5

@@ 2.3 Atomistic Molecular Dynamics simulations (MD)
Classical Molecular Dynamics (MD) simulations were carried out using the Large-scale Atomic/Molecular Massively Parallel Simulator LAMMPS [61]. A cut-off distance of Å is used for both Lennard Jones and Coulomb interactions, and a time-step of fs is utilized. Using a particle-particle particle-mesh solver (PPPM), long-range electrostatic interactions are resolved with an accuracy of 10-4 by mapping atomic charges to a 3D mesh, solving Poisson’s equation on the mesh via 3D Fast Fourier Transforms (FFTs), and interpolating electric fields from mesh points back to atoms. [62, 63]. The electrolyte is prepared using an DMSO:(EMIM+, BF) ratio of 3:2, 1M (Li+, TFSI-), and includes 6000 DMSO molecules, 1857 (EMIM+, BF) pairs , 566 (Li+, TFSI-), 14 ZnI2 and 14 CO2 molecules, giving a total number of atoms of 113142. Each molecule geometry was optimized using the Generalized Amber Force Field (GAFF) [64] and the Universal Force Field force fields (UFF) [65], successively. Then the Automated force field Topology Builder (ATB) is used to obtain the final GROMOS FF parameters for TFSI-, DMSO, EMIM+, BF, and CO2 [66, 67, 68]. We note that the ATB parameter generation considers the following: Lennard-Jones parameters refined against experimental solvation and pure liquid properties, ATB parametrization and validation using a single-range 14 Å cutoff for both Lennard-Jones and Coulomb interactions, atomic charges fitted to quantum mechanical (QM) electrostatic potentials and bonded parameters are assiged using force constants estimated at B3LYP/6-31G* level of theory, resulting in robust symmetry routines for assigning identical parameters to chemically equivalent atoms, bonds, and angles. For a comprehensive understanding of the ATB parameter generation process, refer to the detailed explanations provided in the original research papers [66, 67, 68]. The UFF interaction potential is used for ZnI2. For the Li+, the interaction potential parameters were taken from the Canongia Lopes & Padua (CL&P) force field for ionic liquids [69, 70]. The initial distribution was generated using the Moltemplate package [71]. In this process, the simulated system constituents are placed randomly into an oversized cubic box. Then, the system was energetically minimized to ensure optimal particle distribution. The energy minimization algorithm consists of adjusting particle coordinates iteratively [72]. A finite, unitless stopping tolerance for energy (Estop=10-7) that represents the energy variation between consecutive iterations divided by the energy magnitude, and a stopping tolerance for force (Fstop=10-8 kcal/Å), are chosen. After that, the system is agitated using the Langevin thermostat [73] at T=900 K for t=1 ns, followed by the same process using the Nosé-Hoover thermostat [61, 73]. Then, to reach the desired ambient condition of temperature and pressure (T=300 K, P=1 bar), the simulated systems were equilibrated in the NPT statistical ensemble with isotropic constant pressure control using Berendsen barostat [61, 74]. Firstly, from T=900K to T=300K at a constant pressure of P=500 bar for t=3 ns. Secondly, the pressure decreases from P=500 bar to P=1 bar at a constant temperature of T=300 K over a time period of t=3 ns. In the final NPT equilibration stage, the systems were equilibrated at room temperature and ambient pressure (T=300 K, P=1 bar) for t=10 ns. To investigate the transport properties under conditions similar to experimental settings, we performed simulations in the NVT (constant number of particles, volume, and temperature) statistical ensemble. We employed the Nosé-Hoover thermostat to maintain a constant temperature of 300 K while keeping the volume fixed.
@@ 2.4 Finite element model
A mathematical framework has been developed for modeling the behavior of prismatic Li-CO2 batteries, drawing inspiration from previous studies on Li-O2 and Li-ion batteries [75, 76, 77]. The model integrates key principles such as mass and current conservation, species transport phenomena, and reaction kinetics within the cathode and separator regions, with the aim of providing deeper insights into the complex mechanisms that occur within the cell during operation. The Li-CO2 cell under investigation consists of several components: a lithium metal sheet serving as the negative electrode, a separator, a porous carbon cathode coated with Sb0.67Bi1.33Te3, and an organic electrolyte that permeates the porous structures, as illustrated schematically in Figure 1 (A). The used electrolyte composition is described in section 2.3. To facilitate electron transfer, current collectors are located at the rear of both electrodes. This configuration enables a thorough examination of the electrochemical and transport processes that govern battery cell performance and limitations. The main objective of this mathematical framework is to provide a comprehensive understanding of the intricate interactions among various components and processes within the Li-CO2 battery cell system.
@@ 2.4.1 Electrode Reactions
where and represent the equilibrium potentials for discharge and charge reactions, respectively [17, 18]. The reaction products (Li2CO3 and C) have limited solubility in the DMSO-based liquid electrolyte (0 mol/L for C and 0.0371 mol/L for Li2CO3). When their concentrations exceed these solubility limits, the electrolyte solution becomes supersaturated [78], leading to the deposition of Li2CO3 and C. This is followed by nucleation and growth of these products in the liquid phase within the porous electrode. These deposited products significantly impact the battery’s performance by covering the active surface area, obstructing pathways for reactive species (Li+ and CO2), and inhibiting further reactions inside the cathode, ultimately contributing to the termination of cell discharge.
@@ 2.4.2 Particle transport at the macroscopic level
By introducing the electrode surface
@@ 2.4.3 Equations of battery electrochemical properties
denotes the starting porosity of the cathode. It is important to mention that the effect of deposits (Li2CO3 and C) with low conductivity on was not taken into account, and it was assumed that would not change even as the battery reactions occurred.
@@
The electrochemical reaction at the lithium metal anode, it is given by the general Butler-Volmer equation as follow [75, 76]:
where is exchange current density for anode, is surface or activated overpotential for reaction at anode, and the other parameter are as described above.
The specific area of the electrode/electrolyte interface in Equation 31 is decreased by the morphology and dynamic change in porosity due to /C solid passivation during electrochemical reaction. This effective local surface area per unit volume of electrode can be commonly written by a geometric relation [75, 76]:
@@
The governing equations consist of partial differential equations that include time and space as independent variables. The dependent variables were simulated using COMSOL Multiphysics 6 finite element analysis software to examine their spatial and temporal changes. The battery cell was segmented into three regions: the anode, separator, and porous cathode. An extra fine mesh was selected for enhanced accuracy. To reduce computation time, a variable-step solver was employed with a relative tolerance of 1 10-4. This solver decreases the total number of time steps required for calculation. The parameters used in the simulations are presented in Tables S1, S2, and S3. The applied current density, iapp, was considered positive when directed from the cathode to the external circuit (Figure 1 (A)). The total current density passing through boundary 3 was designated as iapp. The Li+ flux at boundaries 1 and 3 (JLi) and CO2 flux at boundary 2 (J) were set to zero. Additionally, the boundary condition ci = 0 was applied to the relevant boundaries (see Table S2). The parameter c was assumed to be equal to the saturation concentration, c, at boundary 3. The boundary conditions are outlined in Table S2. Discharge simulations were conducted with applied current densities (iapp) ranging from 0.1 to 1 mA/cm2, defined as positive in the direction from the cathode to the external circuit.
@@ 3 Results and discussion
3.1 Electrical conductivity from DFT, AIMD and Kubo-Greenwood formula
Calculated dynamic electrical conductivities for both Sb0.67Bi1.33Te3 and Li2CO3 are shown in Figure 2 (a1) and (b1) for a large frequency range and in Figures 2 (a2) and (b2) for a small frequency range, for different values of E. The static electrical conductivities at zero frequency are calculated by interpolating small frequency data shown in Figures 2 (a2) and (b2). The static electrical conductivities calculated for different values of E show the conductive behavior of Sb0.67Bi1.33Te3 (a conductivity of the order of S/m) and insulative behavior of Li2CO3 (a conductivity of the order of S/m). Our computational analysis reveals a static electrical conductivity of S/m for Sb0.67Bi1.33Te3, indicating its strong electrical conducting properties. This finding aligns with the well-established characteristics of bismuth telluride (Bi2Te3) and its alloys, which are extensively employed in thermoelectric devices due to their good electrical conductivity [85, 86]. The high conductivity can be attributed to the semiconductor nature of Bi2Te3 and the antimony (Sb) doping effect, which augments the carrier concentration [87]. Experimental investigations on Bi2Te3 have reported electrical conductivity values as high as S/m at 323 K [87], further corroborating the computational results. Our calculations reveal that Li2CO3 exhibits a static electrical conductivity of S/m at 300 K, confirming its highly insulating nature. This result closely aligns with experimental measurements of S/m at 393 K [88]. The slight difference between calculated and experimental values is consistent with the expected increase in conductivity at higher temperatures, underscoring the reliability of our computational approach. The insulating nature of Li2CO3 can be attributed to its ionic nature and the strong ionic bonds that hinder electron mobility [89]. In Li2CO3, Li+ ions act as the primary charge carriers rather than electrons, a characteristic typical of ionic compounds. The low conductivity of Li2CO3 presents challenges to battery performance, including high internal resistance and passivation of the electrode surface [90, 91]. To address these challenges, potential solutions include promoting Li2CO3 decomposition through conductive additives or catalysts, designing porous electrodes for increased surface area, or directing the discharge reaction towards products that are easier to decompose (such as Li2C2O4).
@@ 3.1 Electrical conductivity from DFT, AIMD and Kubo-Greenwood formula
To comprehensively understand the formation of Li2CO3 and C on the Sb0.67Bi1.33Te3 catalyst surface, we calculated the Gibbs free energy () changes for various elementary reaction steps on this surface as shown below (steps 1-5) similar to the calculations done in Ref. [4]. Here, we used a slightly bigger supercell (as shown in Figure 2 (a1)) compared to the work done in Ref. [4]. The reaction pathway revealed a sequence of intermediate stages, beginning with the adsorption of CO on a Bi site with a Bi-O distance of 3.5 Å, which was marginally endergonic ( eV). The coupled addition of Li+ and an electron to the adsorbed CO formed the LiCO intermediate, a highly exergonic step ( eV). Subsequent interactions between two LiCO intermediates produced Li2CO and CO, with a of eV. Further, the conversion of Li2CO and CO into Li2CO3 and CO∗ ( eV) marked a pivotal step in the pathway. Finally, the generation of carbon was facilitated through the reaction of CO and CO2, an energetically uphill step ( eV) that also produced CO on the surface. These results, illustrated in Figure 2 (c), highlight the essential role of coupled electron-cation transfer processes in the growth of discharge products. The results in this paper are consistent with the work reported in Ref. [4]. The Gibbs free energy of the reaction was calculated using the following equation:
@@
We considered ZPE and entropy changes in Step-1 and assumed that these changes for other elementary steps are small and therefore negligible. The potential-dependent reaction energies were calculated based on the computational hydrogen electrode (CHE) model with reference to the Li/Li+ electrode [92, 93]. The potential-dependent reaction energies profile is shown in Figure 2 (c), and the respective reaction energies of each step are shown in Table 1.
To level up the electrochemical step, a potential was applied to the formation of LiCO2. As can be seen in Figure 2 (c), this potential occurs at 2.3 V versus Li/Li+ electrode [4].
@@
These insights into the reaction mechanism have significant implications for catalyst design and overall battery performance. The strong affinity of carbon for the catalyst surface, coupled with the energetically unfavorable step of its formation, suggests that catalyst degradation through carbon accumulation could be a concern over long-term battery operation. This highlights the need for strategies to mitigate carbon build-up, such as developing self-cleaning catalysts or implementing periodic regeneration protocols.
Furthermore, the spatial separation of Li2CO3 formation from the immediate catalyst surface underscores the importance of electrolyte engineering in Li-CO2 battery design. Since Li2CO3 formation appears to occur away from the catalyst, the electrolyte’s ability to facilitate Li2CO3 formation, dissolution, and transport becomes crucial. This suggests that future research should focus on developing electrolytes or additives that can enhance Li2CO3 solubility or promote its formation and decomposition.

### v6

@@ 2.3 Atomistic Molecular Dynamics Simulations (MD)
Classical Molecular Dynamics (MD) simulations were performed using LAMMPS [44] to investigate the electrolyte system relevant to Li-CO2 batteries. The simulated electrolyte consisted of DMSO:(EMIM+, BF) at a 3:2 ratio, 1 M (Li+, TFSI-), and included 6000 DMSO molecules, 1857 (EMIM+, BF) pairs, 566 (Li+, TFSI-), 14 ZnI2, and 14 CO2 molecules, totaling 113,142 atoms. Initial configurations were generated and energy minimized, followed by agitation using the Langevin and Nosé-Hoover thermostats. The system was equilibrated to ambient conditions (300 K, 1 bar) in stages using the Berendsen barostat, and transport properties were evaluated under experimental-like conditions in the NVT ensemble at 300 K. Further details on force field parameterization, cutoff schemes, and equilibration protocols are provided in the Supplementary Information.
@@ 2.4 Finite element model
A mathematical framework has been developed for modeling the behavior of prismatic Li-CO2 batteries, drawing inspiration from previous studies on Li-O2 and Li-ion batteries [45, 46, 47]. The model integrates key principles such as mass and current conservation, species transport phenomena, and reaction kinetics within the cathode and separator regions to provide deeper insights into the complex mechanisms that occur within the cell during operation. The Li-CO2 cell under investigation consists of several components: a lithium metal sheet serving as the negative electrode, a separator, a porous carbon cathode coated with Sb0.67Bi1.33Te3, and a liquid electrolyte that permeates the porous structures, as illustrated schematically in Figure 2. The used electrolyte composition is described in section 2.3. To facilitate electron transfer, current collectors are located at the rear of both electrodes. This configuration enables a thorough examination of the electrochemical and transport processes that govern battery cell performance and limitations. The main objective of this mathematical framework is to provide a comprehensive understanding of the intricate interactions among various components and processes within the Li-CO2 battery cell system.
@@ 2.4.1 Electrode Reactions
where and represent the equilibrium potentials for discharge and charge reactions, respectively [17, 18]. The reaction products (Li2CO3 and C) have limited solubility in the DMSO-based liquid electrolyte (0 mol/L for C and 0.0371 mol/L for Li2CO3). When their concentrations exceed these solubility limits, the electrolyte solution becomes supersaturated [48], leading to the deposition of Li2CO3 and C. This is followed by the nucleation and growth of these products in the liquid phase within the porous electrode. These deposited products significantly impact the battery’s performance by covering the active surface area, obstructing pathways for reactive species (Li+ and CO2), and inhibiting further reactions within the cathode, thereby limiting the extent of cell discharge.
@@ 2.4.2 Particle transport at the macroscopic level
By introducing the electrode surface :
@@
| Symbol | Value | Description | Reference |
| Parameters used in the battery assembly | |||
| Lsep | 50 m | Thickness of Separator | This work |
| Lpos | 2 mm | Thickness of positive electrode | This work |
| Ac | 1 cm2 | Cross-sectional area | This work |
| 0.73 | Initial porosity of positive electrode | This work | |
| m | 0.4 mg | mass of active cathode material | This work |
| 0.5 | Porosity of separator | ||
| 1- | Initial active material solid fraction of cathode | This work | |
| cLi,0 | 1000 mol/m3 | Initial concentration of Li+ in electrolyte | This work |
| Parameters calculated using DFT/AIMD simulation | |||
| 1.9 105 S/m | Conductivity of positive electrode | This work | |
| 1.8 10 S/m | electrical conductivity of Li2CO3 film | This work | |
| Parameters calculated using MD simulation | |||
| D | 3.17 10-9 m2/s | Diffusion coefficient of Li+ in electrolyte | This work |
| D | 1.2 10-9 m2/s | CO2 diffusion coefficient | This work |
| 6.77 mS/cm | Conductivity of Li+ in electrolyte | This work | |
| t+ | 0.91 | Transference number of Li+ in electrolyte | This work |
| General parameters | |||
| i0,Li | 0.965 A/m2 | Exchange current density for anode | [51, 46] |
| dlnf±/dlnc | -1.03 | Activity dependence | [51, 46] |
| 2.8 V | Equilibrium potential for discharge | [17] | |
| 3.8 V | Equilibrium potential for charge | [18] | |
| rpos,0 | 25 nm | Particle radius in the positive electrode | [46] |
| a0 | 3/rpos,0 | Initial active specific surface area | [46] |
| n | 4 | Number of transferred electrons | [17] |
| c | 0.0371 mol/L | Solubility limit of Li2CO3 in DMSO | |
| cC,max | 0 mol/L | Solubility limit of Carbon in DMSO | |
| 2140 kg/m3 | Density of Li2CO3 | ||
| M | 73.89 10-3 kg/mol | Molecular weight of Li2CO3 | |
| 2.11 g/cm3 | Density of carbon | ||
| T | 300 K | Temperature | |
| Assumed parameters | |||
| ka | 1.11 10-15 m/s | Anodic reaction rate coefficient | Assumed |
| kc | 3.14 10-15 m7/s/mol2 | Cathodic reaction rate coefficient | Assumed |
| c | 4.8 mol/m3 | Initial CO2 concentration in positive electrode | Assumed |
| Variable | Value | Boundary |
| Total current density | iapp | 3 |
| c | c | 3 |
| 0 V | 1 | |
| 0 | 2 | |
| 0 | 1, 3 |
| Variable | Value | Domain |
| 2.8 V | Porous cathode | |
| 0 V | Separator | |
| 0 V | Porous cathode, Separator | |
| cLi,0 | Porous cathode, Separator | |
| c | Porous cathode | |
| c | Porous cathode | |
| cC,max | Porous cathode | |
| Porous cathode |
@@ 2.4.3 Equations of battery electrochemical properties
denotes the starting porosity of the cathode. It is important to mention that the effect of deposits (Li2CO3 and C) with low conductivity on was not taken into account, and it was assumed that would not change even as the battery reactions occurred. This assumption is primarily justified by the demonstrated structural and chemical stability of the catalyst under cycling conditions [4]. In our continuum model, we further decouple electron transport through the deposited Li2CO3 and carbon byproducts, treating their intrinsic conductivities separately from that of the catalyst. This approach ensures that the catalyst’s role in maintaining electronic percolation pathways is accurately represented. However, it is important to acknowledge the potential limitations of this simplification. If the catalyst were not stable-for example, if it were susceptible to pulverization or surface passivation-its conductivity could decrease during cycling. Such changes might result in localized current hotspots or capacity fade, which would not be captured by the current model. Therefore, while the assumption is reasonable for stable catalyst systems, it may not be universally applicable to all materials or operational conditions.
@@
The electrochemical reaction at the lithium metal anode is given by the general Butler-Volmer equation as follows [45, 46]:
where is the exchange current density for the anode, is the surface or activated overpotential for the reaction at the anode, and the other parameters are as described above.
The specific area of the electrode/electrolyte interface in Equation 31 is decreased by the morphology and dynamic change in porosity due to /C solid passivation during the electrochemical reaction. This effective local surface area per unit volume of the electrode can be commonly written by a geometric relation [45, 46]:
@@
The governing equations consist of partial differential equations that include time and space as independent variables. The dependent variables were simulated using COMSOL Multiphysics 6 Finite Element Analysis software to examine their spatial and temporal changes. The battery cell was segmented into three regions: the anode, separator, and porous cathode. An extra fine mesh was selected for enhanced accuracy. To reduce computation time, a variable-step solver was employed with a relative tolerance of 1 10-4. This solver decreases the total number of time steps required for the calculation. The parameters used in the simulations are presented in Tables 1, 2, and 3. The applied current density, iapp, was considered positive when directed from the cathode to the external circuit (Figure 2). The total current density passing through boundary 3 was designated as iapp. The Li+ flux at boundaries 1 and 3 (JLi) and CO2 flux at boundary 2 (J) were set to zero. Additionally, the boundary condition ci = 0 was applied to the relevant boundaries (see Table 2). The parameter c was assumed to be equal to the saturation concentration, c, at boundary 3. The boundary conditions are outlined in Table 2. Discharge simulations were conducted with applied current densities (iapp) ranging from 0.1 to 1 mA/cm2, defined as positive in the direction from the cathode to the external circuit.
@@ 3 Results and discussion
3.1 Electrical conductivity from DFT, AIMD, and Kubo-Greenwood formula
The calculated dynamic electrical conductivities for Sb0.67Bi1.33Te3 and Li2CO3 are shown in Figure 3 (a1) and (b1) across a broad frequency range, and in Figures 3 (a2) and (b2), which focus on low frequencies to enable extrapolation to 0 Hz, with E values ranging from 0.22 eV to 0.46 eV. Static electrical conductivities at zero frequency were determined by extrapolating the low-frequency data in Figures 3 (a2) and (b2). The calculated conductivities are within the experimental order of magnitude. E values between 0.34 eV and 0.42 eV were selected because they show convergence with experimental conductivity data within this range. Furthermore, varying the electrical conductivity within the computational uncertainty does not significantly affect the predicted battery properties in the FEA. The static electrical conductivities calculated for different values of E show the conductive behavior of Sb0.67Bi1.33Te3 (a conductivity of the order of S/m) and insulative behavior of Li2CO3 (a conductivity of the order of S/m). Our computational analysis reveals a static electrical conductivity of S/m for Sb0.67Bi1.33Te3, indicating its strong electrical conducting properties. This finding aligns with the well-established characteristics of bismuth telluride (Bi2Te3) and its alloys, which are extensively employed in thermoelectric devices due to their good electrical conductivity [56, 57]. The high conductivity can be attributed to the semiconductor nature of Bi2Te3 and the antimony (Sb) doping effect, which augments the carrier concentration [58]. Experimental investigations on Bi2Te3 have reported electrical conductivity values as high as S/m at 323 K [58], further corroborating the computational results. Our calculations reveal that Li2CO3 exhibits a static electrical conductivity of S/m at 300 K, confirming its highly insulating nature. This result closely aligns with experimental measurements of S/m at 393 K [59]. The slight difference between calculated and experimental values is consistent with the expected increase in conductivity at higher temperatures, underscoring the reliability of our computational approach. The insulating nature of Li2CO3 can be attributed to its ionic nature and the strong ionic bonds that hinder electron mobility [60]. In Li2CO3, Li+ ions act as the primary charge carriers rather than electrons, a characteristic typical of ionic compounds. The low conductivity of Li2CO3 presents challenges to battery performance, including high internal resistance and passivation of the electrode surface [61, 62].
The calculated electrical conductivities are used in the continuum model to characterize electron percolation in the cathode and the deposited Li2CO3 during the battery charge/discharge cycle.
@@ 3.1 Electrical conductivity from DFT, AIMD, and Kubo-Greenwood formula
To comprehensively understand the formation of Li2CO3 and C on the Sb0.67Bi1.33Te3 catalyst surface, we calculated the Gibbs free energy () changes for various elementary reaction steps on this surface as shown below (steps 1-5) similar to the calculations done in Ref. [4]. Here, we used a slightly bigger supercell (as shown in Figure 3 (a1)) compared to the work done in Ref. [4]. The reaction pathway revealed a sequence of intermediate stages, beginning with the adsorption of CO on a Bi site with a Bi-O distance of 3.5 Å, which was marginally endergonic ( eV). The coupled addition of Li+ and an electron to the adsorbed CO formed the LiCO intermediate, a highly exergonic step ( eV). Subsequent interactions between two LiCO intermediates produced Li2CO and CO, with a of eV. Further, the conversion of Li2CO and CO into Li2CO3 and CO∗ ( eV) marked a pivotal step in the pathway. Finally, the generation of carbon was facilitated through the reaction of CO and CO2, an energetically uphill step ( eV) that also produced CO on the surface. These results, illustrated in Figure 4, highlight the essential role of coupled electron-cation transfer processes in the growth of discharge products. The results in this paper are consistent with the work reported in Ref. [4]. The Gibbs free energy of the reaction was calculated using the following equation:
@@
In the calculation of Gibbs free energy profiles, full corrections including zero-point energy (ZPE) and entropy were applied explicitly to the initial gas-to-surface adsorption step of CO2, where such contributions are known to be significant due to the loss of translational and rotational entropy. For subsequent steps involving transformations between adsorbed intermediates (e.g., Li2CO3, Li2CO2, Li2CO3, and C–CO3 species), we assumed that ZPE and entropy effects largely cancel out, as these involve similar bonding environments and restricted surface-bound species. This approach is widely used in heterogeneous catalysis modeling and has been shown to introduce minimal quantitative error, while preserving the qualitative accuracy of energy landscapes and mechanistic conclusions [63]. Accordingly, our analysis emphasizes the dominant electronic energy contributions for these surface processes, which govern the reaction thermodynamics and barrier-lowering effects central to catalyst performance. The potential-dependent reaction energies were calculated based on the computational hydrogen electrode (CHE) model with reference to the Li/Li+ electrode [64, 65]. The potential-dependent reaction energies profile is shown in Figure 4, and the respective reaction energies of each step are shown in Table 4.
To level up the electrochemical step, a potential was applied to the formation of LiCO2. As can be seen in Figure 4, this potential occurs at 2.3 V versus Li/Li+ electrode [4].
@@
These insights into the reaction mechanism have significant implications for catalyst design and overall battery performance. The strong affinity of carbon for the catalyst surface, coupled with the energetically unfavorable step of formation, suggests that catalyst degradation through carbon accumulation could be a concern over long-term battery operation. This highlights the need for strategies to mitigate carbon build-up, such as developing self-cleaning catalysts or implementing periodic regeneration protocols.
Furthermore, the spatial separation of Li2CO3 formation from the immediate catalyst surface underscores the importance of electrolyte engineering in Li-CO2 battery design. Since Li2CO3 formation appears to occur away from the catalyst, the electrolyte’s ability to facilitate Li2CO3 formation, dissolution, and transport becomes crucial.
The continuum model considers the formation of Li2CO3 and carbon in the cathode. We assume that the catalyst increases the reaction rate of Li2CO3/C formation and use a reaction rate constant on the order of 10-15 m7 s-1 mol-2. This value is higher than the typical value assumed for Li2O2 formation in Li-O2 batteries with non-catalyzed carbon cathodes, which is on the order of 10-17 m7 s-1 mol-2 [46].